OpenStudy (phebe):

HELPSSSSS FOR MEDALSSSS 17. Simplify by rationalizing the denominator: (-√18)/√12

OpenStudy (anonymous):

didnt i solve this yesterday

OpenStudy (anonymous):

so what seems to be the problem?

OpenStudy (phebe):

i forgot wat u told me the answer was @aajugdar

OpenStudy (anonymous):

oh wait wait,i did some mistake there

OpenStudy (phebe):

ok

OpenStudy (anonymous):

when you get \[-3\sqrt{2}/2\sqrt{3} = -\sqrt{3}/\sqrt{2}\]

OpenStudy (anonymous):

that will be the final answer

OpenStudy (anonymous):

shamim's method is correct,i didnt rationalise the denominator my bad

OpenStudy (phebe):

ok thank tou very much

OpenStudy (anonymous):

so if you multiply numarator and denominator by root 2 u will get \[-\sqrt{6}/\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

rationalising ~ removing root from number

OpenStudy (anonymous):

so yes \[-\sqrt{6}/2 \] is correct

OpenStudy (phebe):

wait is that the answer

OpenStudy (anonymous):

yep

OpenStudy (phebe):

ok thanx